OpenStudy (anonymous):

How do I express this equation into simplest radical form? x square root of 2 = square root of 14. I'm confused, someone please tell me the steps! D:

OpenStudy (anonymous):

x\[\sqrt{2} = \sqrt{14}\]

OpenStudy (anonymous):

except the x is beside square root of 2

OpenStudy (anonymous):

\[x \sqrt{2}=\sqrt{14}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

the answer is square root of 7.. o.o but why and how?

OpenStudy (anonymous):

Divide both sides by \[\sqrt{2}\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

\[x = \frac{\sqrt{14}}{ \sqrt{2}}\]

OpenStudy (anonymous):

and that's it?

OpenStudy (anonymous):

\[\frac{\sqrt[n]{a}}{\sqrt[n]{b}} = \sqrt[n](\frac{a}{b})\]

OpenStudy (anonymous):

So you can divide 14 by 2 and keep it under 1 root

OpenStudy (anonymous):

ohh.. thank you!

OpenStudy (anonymous):

np